# On the Generalization of SFT: A Reinforcement Learning Perspective with Reward Rectification

### v2

@@ Abstract
In this work, we present a simple yet theoretically motivated improvement to Supervised Fine-Tuning (SFT) for the Large Language Model (LLM), addressing its limited generalization compared to reinforcement learning (RL). Through mathematical analysis, we reveal that standard SFT gradients implicitly encode a problematic reward structure that may severely restrict the generalization capabilities of model. To rectify this, we propose Dynamic Fine-Tuning (DFT), stabilizing gradient updates for each token by dynamically rescaling the objective function with the probability of this token. With just a single-line change, the method outperforms standard SFT on multiple difficult benchmarks and base models, from math reasoning to code generation and multi-modal tasks, demonstrating improved generalization. Additionally, DFT achieves competitive results in offline RL settings, providing an effective yet streamlined alternative. By bridging theoretical insights with practical solutions, this work advances the state of SFT. The code will be available at https://github.com/yongliang-wu/DFT.
@@ 1 Introduction
Supervised Fine-Tuning (SFT), which adapts models to expert demonstrations, has become the standard post-training paradigm for Large Language Models (LLMs). It enables efficient task adaptation and capability enhancement (Chung et al., 2024; Zhang et al., 2024b; Sanh et al., 2022; Ouyang et al., 2022), and is popular for its ease of implementation and rapid acquisition of expert-like behaviors (Wei et al., 2022; Zhou et al., 2023). Despite these advantages, SFT often shows limited generalization compared to reinforcement learning (RL) (Chu et al., 2024; Ouyang et al., 2022; Christiano et al., 2017; Bai et al., 2022; Huan et al., 2025; Swamy et al., 2025). RL leverages explicit reward or verification signals to explore diverse strategies and thus generalizes better. However, RL requires substantial computation, careful hyperparameter tuning, and explicit reward signals—conditions often impractical in real-world settings (Schulman et al., 2017; Ouyang et al., 2022; Sheng et al., 2025; Strubell et al., 2019; Liu & Yin, 2024; Winsta, 2025). Moreover, RL can struggle to recover expert-like behaviors that SFT captures efficiently (Mandlekar et al., 2022; Chen et al., 2025b).
@@
In this work, we address this gap with a mathematical analysis of the connection between SFT and RL. We show that SFT’s gradient update is equivalent to a policy gradient with a specific, implicitly defined reward. Crucially, this reward is (i) sparse, and (ii) inversely proportional to the model’s probability of expert actions (see equation 6). As a result, when the model assigns low probability to expert actions, the gradient becomes excessively large—yielding an ill-posed reward structure and unstable, overfitting-prone optimization.
Building on this insight, we propose Dynamic Fine-Tuning (DFT), a principled fix. Our method rescales the SFT objective at each token by its probability, canceling the distortion introduced by inverse-probability weighting. This transforms the SFT gradient from an unstable, biased estimator into a stable, uniformly weighted update rule.
Empirically, DFT delivers substantial improvements. On the Qwen-2.5-Math series (Qwen Team et al., 2024b) fine-tuned with NuminaMath-CoT (LI et al., 2024), DFT yields gains several times larger than standard SFT. More importantly, unlike SFT, which often degrades on challenging benchmarks such as OlympiadBench (He et al., 2024), AIME 2024 (American Institute of Mathematics, 2024), and AMC 2023 (Mathematical Association of America, 2023), our method consistently improves performance and generalization. These improvements hold across models, scales, and data sizes (Table 1, Figure 1), and extend to code generation and multimodal reasoning (Tables 3, 4).
We further test DFT in off-policy RL settings (Table 2), where dense rewards are available (Levine et al., 2020). Our method not only outperforms offline RL approaches such as DPO (Rafailov et al., 2023) and RAFT (Dong et al., 2023; Ahn et al., 2024), but also achieves competitive or superior performance to online methods like GRPO and PPO on math tasks with Qwen2.5-Math-1.5B. Unlike these RL methods, DFT requires neither a reference model nor large batch sizes, making it a simpler and more resource-efficient alternative.
To understand its effect, we analyze token probability distributions after training (Figure 2). While traditional SFT uniformly pushes probabilities toward the training set, DFT selectively increases some while reducing others. In particular, the proportion of less strongly fitted tokens rises, suggesting improved regularization and reduced overfitting. We provide further discussion in Appendix A.4.
@@ 2 Related Work
The trade-off between supervised fine-tuning (SFT) and reinforcement learning (RL) is central to the alignment of large language models. SFT is widely adopted due to its simplicity and efficiency in imitating expert demonstrations (Chung et al., 2024; Zhou et al., 2023; Wei et al., 2022), analogous to behavioral cloning in robotics (Sammut, 2011; Mandlekar et al., 2022). However, the literature consistently highlights its limitations, particularly the tendency to overfit and generalize poorly compared to RL, which leverages reward signals to discover more robust policies (Ouyang et al., 2022; Christiano et al., 2017; Bai et al., 2022; Swamy et al., 2025; Zhang et al., 2025). A recent systematic comparison by Chu et al. (2024) across textual and visual domains confirms this distinction, concisely summarized as “SFT memorizes while RL generalizes.” They further show that SFT remains indispensable as an initialization step, stabilizing output formatting prior to effective RL training. Nonetheless, RL faces significant practical hurdles, including computational expense, sensitivity to hyperparameters, and the requirement of an explicit reward function, all of which constrain its applicability (Schulman et al., 2017; Strubell et al., 2019; Sheng et al., 2025).
@@
To combine the strengths of both paradigms, much recent work has pursued hybrid approaches. The most common strategy involves SFT pretraining followed by RL-based refinement with a learned reward model, as popularized by InstructGPT (Ouyang et al., 2022). More recent methods interleave SFT and RL updates to improve stability and performance (Sheng et al., 2025; Liu et al., 2025; Qiu et al., 2025). Other approaches, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023), bypass reward modeling entirely by directly optimizing policies on preference data, thereby unifying imitation and reinforcement signals within a single loss function. Chen et al. (2025a) introduce Negative-aware Fine-Tuning (NFT), which models incorrect generations via an implicit negative policy, enabling self-improvement without explicit feedback. While powerful, these methods rely on reward signals, preference pairs, or negative samples. They enrich the training pipeline but do not fundamentally improve SFT in its native setting, where only positive demonstrations are available. Our work instead focuses on enhancing SFT itself without requiring external feedback.
@@
A complementary line of theoretical research seeks to unify SFT and RL under a common formalism. Du et al. (2025) reinterpret RLHF as a reward-weighted variant of SFT, preserving reliance on an explicit reward. Wang et al. (2025) show that SFT can be cast as RL with an implicit reward, proposing adjustments such as smaller learning rates to manage the vanishing KL constraint. Abdolmaleki et al. (2025) analyze learning from both positive and negative feedback, studying how their balance affects convergence. Qin & Springenberg (2025) view SFT as a lower bound of RL and introduce importance weighting based on the data-generating policy. While these works establish connections between SFT and RL through weighting, they do not provide a precise mathematical equivalence between the SFT gradient and the offline policy gradient. Some methods approximate this connection in practice by reweighting training losses. For instance, MixCE (Zhang et al., 2023) combines the forward and reverse KL divergences to form a unified objective, while GOLD (Pang & He, 2021) adopts offline RL with demonstrations, introducing reliance on an unknown demonstration distribution $\pi_{b}$ and a restrictive $1/N$ assumption. In contrast, our work rigorously establishes this equivalence, pinpointing the inverse-probability weighting term in SFT as the key distinction. This observation directly motivates our proposed solution: multiplying the SFT loss by the model probability to eliminate the weighting.
@@ 3.2 Unify SFT and RL Gradient Expression
This form of the SFT gradient now aligns closely with the policy gradient equation 4. Thus we can see, conventional SFT is precisely an on-policy-gradient with the reward as an indicator function of matching the expert trajectory but biased by an importance weighting $1/\pi_{\theta}$. More detailed derivations can be found in Appendix A.3.
Given the unavoidable sparsity of reward signals in the SFT setting, we identify the importance sampling weight $1/\pi_{\theta}$ as a fundamental cause of the poor generalization of SFT relative to RL. When the model assigns low probability to the expert response, the weight $w$ becomes excessively large, leading to an ill-posed reward structure. In such cases, the gradient can grow disproportionately, making the training unstable and prone to overfitting. This issue is further exacerbated by the extreme sparsity of the reward function, since $r(x,y)=\mathbf{1}[y=y^{\star}]$ is non-zero only when the model exactly matches the expert output. As a result, optimization tends to overfit rare exact-match demonstrations, undermining the model’s ability to generalize beyond training data.
@@ 3.3 Proposed Method
$\nabla_{\theta}\mathcal{L}_{\mathrm{DFT}}(\theta)=\nabla_{\theta}\mathcal{L}_{\mathrm{SFT}}(\theta)\;\cdot\;\operatorname{sg}(\frac{1}{w})=\nabla_{\theta}\mathcal{L}_{\mathrm{SFT}}(\theta)\;\cdot\;\operatorname{sg}(\pi_{\theta}(y^{\star}\mid x)).$ (7)
@@
$\mathcal{L}_{\text{DFT}}(\theta)=\mathbb{E}_{(x,y^{\star})\sim\mathcal{D}}\Bigl[\operatorname{sg}\big(\pi_{\theta}(y^{\star}\mid x)\big)\log\pi_{\theta}(y^{\star}\mid x)\Bigr].$ (8)
@@ 4.1 Main Experiment - Mathematical Reasoning Task
4.1.1 Setup and Implementation Details
To efficiently manage computational resources, We andomly sample 100,000 instances from the the NuminaMath-CoT dataset (LI et al., 2024) for training. We conduct experiments using multiple models, including Qwen2.5-Math-1.5B, Qwen2.5-Math-7B (Qwen Team et al., 2024a), LLaMA-3.2-3B, LLaMA-3.1-8B (Dubey et al., 2024), and DeepSeekMath-7B (Shao et al., 2024).
Our implementation builds upon the verl framework (Sheng et al., 2025), using recommended SFT hyper-parameters. Specifically, we employ the AdamW optimizer with learning rates of $5\times 10^{-5}$ for all models except the LLaMA-3.1-8B-Base, for which we adopt a lower learning rate of $2\times 10^{-5}$. We set the mini-batch size to 256 and the maximum input length to 2048 tokens. The learning rate follows a cosine decay schedule with a warm-up ratio of 0.1. We evaluate on benchmarks including Math500 (Hendrycks et al., 2021), Minerva Math (Lewkowycz et al., 2022), Olympiad Bench (He et al., 2024), AIME 2024 (American Institute of Mathematics, 2024), and AMC 2023 (Mathematical Association of America, 2023) through the official Qwen2.5-Math evaluation pipeline (Qwen Team et al., 2024a). Each model uses the default chat template and Chain-of-Thought (CoT) prompting to stimulate step-by-step reasoning. All reported results represent average accuracy across 16 decoding runs, evaluated with a temperature of 1.0 and maximum generation length of 4096 tokens.
@@ 4.1.2 Main Results
We also report the results of parameter-efficient fine-tuning (PEFT) training setting (Hu et al., 2022) and training on the OpenR1-Math dataset (Hugging Face, 2025) with better quality in Appendix A.7 and Appendix A.6, respectively. Comparison and Discussion with the concurrent method iw-SFT (Qin & Springenberg, 2025) is provided in Appendix A.5.
@@ 4.2 Exploratory Experiment - Offline RL Setting
4.2.1 Setup and Implementation Details
We sample responses for 100,000 math questions using a temperature of 1.0 and generate four responses per question from the base model itself. Correct responses are identified using math verify and retained as training data, resulting in approximately 140,000 examples. For DPO training, we construct 100,000 positive–negative preference pairs from the generated responses.
We compare DFT with representative offline RL methods, including DPO (Rafailov et al., 2023) and RFT (Dong et al., 2023; Ahn et al., 2024), as well as online RL methods PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024). For RFT and DFT, the training setup follows the configuration in Section 4.1. For DPO, we use the ms-swift (Zhao et al., 2024) with a learning rate of $1\times 10^{-6}$, batch size of 128, and a warmup ratio of 0.05. For PPO and GRPO, training is performed using the verl (Sheng et al., 2025) with a learning rate of $1\times 10^{-6}$, batch size of 256, and a warmup ratio of 0.1. We set the number of response $n=4$ for GRPO.
4.2.2 Results

### v3

# arXiv:2508.05629v3[cs.LG] 27 Feb 2026

@@ Abstract
In this work, we present a simple yet theoretically motivated improvement to Supervised Fine-Tuning (SFT) for the Large Language Model (LLM), addressing its limited generalization compared to reinforcement learning (RL). Through mathematical analysis, we reveal that standard SFT gradients implicitly encode a problematic reward structure that may severely restrict the generalization capabilities of model compared to RL. To rectify this, we propose Dynamic Fine-Tuning (DFT), stabilizing gradient updates for each token by dynamically rescaling the objective function with the probability of this token. With just a single-line change, the method outperforms standard SFT on multiple difficult benchmarks and base models, from math reasoning to code generation and multi-modal tasks, demonstrating improved generalization. Additionally, DFT achieves competitive results in offline RL settings, providing an effective yet streamlined alternative. By bridging theoretical insights with practical solutions, this work advances the state of SFT. The source code will be available at https://github.com/yongliang-wu/DFT.
@@ 1 Introduction
Supervised Fine-Tuning (SFT), which adapts models to expert demonstrations, has become the standard post-training paradigm for Large Language Models (LLMs) (Zhang et al., 2025c; 2024a; b; a; Wang et al., 2025b). It enables efficient task adaptation and capability enhancement (Chung et al., 2024; Zhang et al., 2024c; Sanh et al., 2022; Ouyang et al., 2022; Chen et al., 2024; 2025a; Fang et al., 2025), and is popular for its ease of implementation and rapid acquisition of expert-like behaviors (Wei et al., 2022; Zhou et al., 2023). Despite these advantages, SFT often shows limited generalization compared to reinforcement learning (RL) (Chu et al., 2024; Ouyang et al., 2022; Christiano et al., 2017; Bai et al., 2022; Huan et al., 2025; Swamy et al., 2025). RL leverages explicit reward or verification signals to explore diverse strategies and thus generalizes better. However, RL requires substantial computation, careful hyperparameter tuning, and explicit reward signals—conditions often impractical in real-world settings (Schulman et al., 2017; Ouyang et al., 2022; Sheng et al., 2025; Strubell et al., 2019; Liu and Yin, 2024; Winsta, 2025). Moreover, RL can struggle to recover expert-like behaviors that SFT captures efficiently (Mandlekar et al., 2022; Chen et al., 2025c).
@@
In this work, we address this gap with a mathematical analysis of the connection between SFT and RL. We show that the gradient update in SFT can be interpreted as a form of policy gradient with a specific, implicitly defined reward under certain assumptions. Crucially, this reward is (i) sparse, and (ii) inversely proportional to the model’s probability of expert actions (see equation 6). As a result, when the model assigns low probability to expert actions, the gradient becomes excessively large, yielding an ill-posed reward structure and unstable optimization (Pascanu et al., 2013; Yang et al., 2019).
Building on this insight, we propose Dynamic Fine-Tuning (DFT), a principled fix. Our method rescales the SFT objective at each token by its probability, canceling the distortion introduced by inverse-probability weighting. This reframing turns the SFT gradient from a potentially unstable and biased estimator into a more stable, more uniformly weighted update rule that behaves closer to an RL-style.
Empirically, DFT delivers substantial improvements. On the Qwen-2.5-Math series (Qwen Team et al., 2024a) fine-tuned with NuminaMath-CoT (LI et al., 2024), DFT yields gains several times larger than standard SFT. More importantly, unlike SFT, which often degrades on challenging benchmarks such as OlympiadBench (He et al., 2024), AIME 2024 (American Institute of Mathematics, 2024), and AMC 2023 (Mathematical Association of America, 2023), our method consistently improves performance and generalization. These improvements hold across models, scales, and data sizes (Table 1, Figure 1), and extend to code generation and multimodal reasoning (Tables 3, 4(Zhao et al., 2025d; Luo et al., 2025; Li et al., 2025a; b) (Zhu et al., 2024b; c; a; 2020).
We further test DFT in off-policy RL settings (Table 2), where dense rewards are available (Levine et al., 2020). Our method not only outperforms offline RL approaches such as DPO (Rafailov et al., 2023) and RAFT (Dong et al., 2023; Ahn et al., 2024), but also achieves competitive or superior performance to online methods like GRPO and PPO on math tasks with Qwen2.5-Math-1.5B (He et al., 2025; Tan et al., 2026; Zhao et al., 2025b; a). Unlike these RL methods, DFT requires neither a reference model nor large batch sizes, making it a simpler and more resource-efficient alternative.
To understand its effect, we analyze token probability distributions after training (Figure 2). While traditional SFT uniformly pushes probabilities toward the training set, DFT selectively increases some while reducing others. In particular, the proportion of less strongly fitted tokens rises, suggesting improved regularization. We provide further discussion in Appendix A.3.
@@ 2 Related Work
The trade-off between supervised fine-tuning (SFT) and reinforcement learning (RL) is central to the alignment of large language models (Song et al., 2024; Chai et al., 2024; Song et al., 2025c; a; Xu et al., 2025; Song et al., 2025b; Zhu et al., 2025b; c; a). SFT is widely adopted due to its simplicity and efficiency in imitating expert demonstrations (Chung et al., 2024; Zhou et al., 2023; Wei et al., 2022), analogous to behavioral cloning in robotics (Sammut, 2011; Mandlekar et al., 2022). However, the literature consistently highlights its limitations, particularly the tendency to overfit and generalize poorly compared to RL, which leverages reward signals to discover more robust policies (Ouyang et al., 2022; Christiano et al., 2017; Bai et al., 2022; Swamy et al., 2025; Zhang et al., 2025d). A recent systematic comparison by Chu et al. (2024) across textual and visual domains confirms this distinction, concisely summarized as “SFT memorizes while RL generalizes.” They further show that SFT remains indispensable as an initialization step, stabilizing output formatting prior to effective RL training. Nonetheless, RL faces significant practical hurdles, including computational expense, sensitivity to hyperparameters, and the requirement of an explicit reward function, all of which constrain its applicability (Schulman et al., 2017; Strubell et al., 2019; Sheng et al., 2025).
@@
To combine the strengths of both paradigms, much recent work has pursued hybrid approaches. The most common strategy involves SFT pretraining followed by RL-based refinement with a learned reward model, as popularized by InstructGPT (Ouyang et al., 2022). More recent methods interleave SFT and RL updates to improve stability and performance (Sheng et al., 2025; Liu et al., 2025; Qiu et al., 2025). Other approaches, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023), bypass reward modeling entirely by directly optimizing policies on preference data, thereby unifying imitation and reinforcement signals within a single loss function. Chen et al. (2025b) introduce Negative-aware Fine-Tuning (NFT), which models incorrect generations via an implicit negative policy, enabling self-improvement without explicit feedback. While powerful, these methods rely on reward signals, preference pairs, or negative samples. They enrich the training pipeline but do not fundamentally improve SFT in its native setting, where only positive demonstrations are available. Our work instead focuses on enhancing SFT itself without requiring external feedback.
@@
A complementary line of theoretical research seeks to unify SFT and RL under a common formalism. Du and others (2025) reinterpret RLHF as a reward-weighted variant of SFT, preserving reliance on an explicit reward. Wang et al. (2025a) show that SFT can be cast as RL with an implicit reward, proposing adjustments such as smaller learning rates to manage the vanishing KL constraint. Abdolmaleki et al. (2025) analyze learning from both positive and negative feedback, studying how their balance affects convergence. Qin and Springenberg (2025) view SFT as a lower bound of RL and introduce importance weighting based on the data-generating policy. While these works establish connections between SFT and RL through weighting, they do not provide a precise mathematical equivalence between the SFT gradient and the offline policy gradient. Some methods approximate this connection in practice by reweighting training losses. For instance, MixCE (Zhang et al., 2023) combines the forward and reverse KL divergences to form a unified objective, while GOLD (Pang and He, 2021) adopts offline RL with demonstrations, introducing reliance on an unknown demonstration distribution $\pi_{b}$ and a restrictive $1/N$ assumption. Kantharaju and Sankar (2022) also provide a clear and insightful exposition of GOLD’s motivation and mechanics from an alternative perspective, offering useful intuition for understanding its underlying design. Zhao et al. (2025c) provide a promising method to combine RL and SFT during training. In contrast, our work offers a more formal perspective on this connection, highlighting the role of the inverse-probability weighting term in shaping the difference between SFT and RL-like updates. This perspective motivates a simple adjustment: multiplying the loss by the model’s token probability to neutralize the weighting.
@@ 3.2 Unify SFT and RL Gradient Expression
This form of the SFT gradient closely resembles the policy gradient in Equation equation 4. Under this formulation, conventional SFT can be interpreted as an on-policy gradient method, where the reward is a sparse indicator function matching the expert trajectory, but biased by an importance weighting term $1/\pi_{\theta}$. We emphasize that this RL-style characterization serves solely as a theoretical lens: both the analysis and subsequent modifications are developed within the RL framework, while the final method remains fully implementable in standard SFT form for computational efficiency. Detailed derivations are provided in Appendix A.2.
Due to the inherently sparse reward signal in the SFT setting, we identify the importance weight $1/\pi_{\theta}$ as a key contributor to SFT’s generalization limitations compared to RL. When the model assigns low probability to the expert response, the resulting weight becomes excessively large, introducing an ill-posed reward landscape. This leads to disproportionately large gradients and training instability. The issue is compounded by the fact that the reward function $r(x,y)=\mathbf{1}[y=y^{\star}]$ is non-zero only for exact matches to the expert outputm causing optimization to overfit rare exact-match samples and weakening the model’s ability to generalize beyond the training data.
@@ 3.3 Proposed Method
$\nabla_{\theta}\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{x\sim\mathcal{D}_{x},\;y\sim\pi_{\theta}(\cdot\mid x)}\bigl[\operatorname{sg}(\frac{1}{w})\cdot{\color[rgb]{0,0,1}w(y\mid x)}\,\nabla_{\theta}\log\pi_{\theta}(y\mid x)\,{\color[rgb]{0,0,1}r(x,y)}\bigr].$ (7)
@@
$\mathcal{L}_{\text{DFT}}(\theta)=\mathbb{E}_{(x,y^{\star})\sim\mathcal{D}}\Bigl[-\operatorname{sg}\big(\pi_{\theta}(y^{\star}\mid x)\big)\log\pi_{\theta}(y^{\star}\mid x)\Bigr].$ (8)
@@ 4.1 Main Experiment - Mathematical Reasoning Task
Implementation details.
To efficiently manage computational resources, We andomly sample 100,000 instances from the the NuminaMath-CoT dataset (LI et al., 2024) for training. We conduct experiments using multiple models, including Qwen2.5-Math-1.5B, Qwen2.5-Math-7B (Qwen Team et al., 2024b), LLaMA-3.2-3B, LLaMA-3.1-8B (Dubey et al., 2024), and DeepSeekMath-7B (Shao et al., 2024). Our implementation builds upon the verl framework (Sheng et al., 2025), using recommended SFT hyper-parameters. Specifically, we employ the AdamW optimizer with learning rates of $5\times 10^{-5}$ for all models except the LLaMA-3.1-8B-Base, for which we adopt a lower learning rate of $2\times 10^{-5}$. We set the mini-batch size to 256 and the maximum input length to 2048 tokens. The learning rate follows a cosine decay schedule with a warm-up ratio of 0.1. We evaluate on benchmarks including Math500 (Hendrycks et al., 2021), Minerva Math (Lewkowycz et al., 2022), Olympiad Bench (He et al., 2024), AIME 2024 (American Institute of Mathematics, 2024), and AMC 2023 (Mathematical Association of America, 2023) through the official Qwen2.5-Math evaluation pipeline (Qwen Team et al., 2024b). Each model uses the default chat template and Chain-of-Thought (CoT) prompting to stimulate step-by-step reasoning. All reported results represent average accuracy across 16 decoding runs, evaluated with a temperature of 1.0 and maximum generation length of 4096 tokens.
@@
We also report the results of parameter-efficient fine-tuning (PEFT) training setting (Hu et al., 2022) and training on the OpenR1-Math dataset (Hugging Face, 2025) with better quality in Appendix A.6 and Appendix A.5, respectively. Comparison and Discussion with the concurrent method iw-SFT (Qin and Springenberg, 2025) is provided in Appendix A.4.
@@ 4.2 Exploratory Experiment - Offline RL Setting
Implementation details.
We sample responses for 100,000 math questions using a temperature of 1.0 and generate four responses per question from the base model itself. Correct responses are identified using math verify and retained as training data, resulting in approximately 140,000 examples. For DPO training, we construct 100,000 positive–negative preference pairs from the generated responses. We compare DFT with representative offline RL methods, including DPO (Rafailov et al., 2023) and RFT (Dong et al., 2023; Ahn et al., 2024), as well as online RL methods PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024). For RFT and DFT, the training setup follows the configuration in Section 4.1. For DPO, we use the ms-swift (Zhao et al., 2024) with a learning rate of $1\times 10^{-6}$, batch size of 128, and a warmup ratio of 0.05. For PPO and GRPO, training is performed using the verl (Sheng et al., 2025) with a learning rate of $1\times 10^{-6}$, batch size of 256, and a warmup ratio of 0.1. We set the number of response $n=4$ for GRPO.